How much money must you deposit into a savings account at the end of each year at 4% interest compounded annually in order to earn $9,778.08 interest during a 20-year period?


I got Future value = 8209.13 + 9778.08 = $17987.21

But the formula to insert is {[(1 + i)^n - 1]/i}.

I keep missing the mark but not sure what I am doing wrong
Answer #1

Answer:

Given

Interest rate i=4%

Number of years n=20

Total interest earned R=$9778.08

Let A be the amount deposited at end of each year.

So Future value of annual Deposit FV=A*((1+r)^n-1)/r

FV=A*((1+4%)^20-1)/4%=29.77808A

So total Interest earned R= FV-20*A=29.77808A-20A=9.77808A

9.77808A=9778.08

A=$1000